What is the pattern?
2 - 1
3 - 3
4 - 6
5 - 10
6 - 15
7 - 21
8 - 28
9 - 36
10- 45
I have no idea what the pattern is
thanks for helping

Answer by stanbon(75887)

What is the pattern?
2 - 1
3 - 3
4 - 6
5 - 10
6 - 15
7 - 21
8 - 28
9 - 36
10- 45
-----------
When n = 2, a(2) = 1
when n = 3, a(3) = a(2)+2 = 3
when n = 4, a(4) = a(3)+3 = 6
when n = 5, a(5) = a(4)+4 = 10
...
...
when n= k; a(k) = a(k-1)+(k-1)
==================================
Cheers,
Stan H.